# State-Space Modeling in Long Sequence Processing: A Survey on Recurrence in the Transformer Era

### v1

@@ Abstract
Effectively learning from sequential data is a longstanding goal of Artificial Intelligence, especially in the case of long sequences. From the dawn of Machine Learning, several researchers engaged in the search of algorithms and architectures capable of processing sequences of patterns, retaining information about the past inputs while still leveraging the upcoming data, without losing precious long-term dependencies and correlations. While such an ultimate goal is inspired by the human hallmark of continuous real-time processing of sensory information, several solutions simplified the learning paradigm by artificially limiting the processed context or dealing with sequences of limited length, given in advance. These solutions were further emphasized by the large ubiquity of Transformers, that have initially shaded the role of Recurrent Neural Nets. However, recurrent networks are facing a strong recent revival due to the growing popularity of (deep) State-Space models and novel instances of large-context Transformers, which are both based on recurrent computations to go beyond several limits of currently ubiquitous technologies. In fact, the fast development of Large Language Models enhanced the interest in efficient solutions to process data over time. This survey provides an in-depth summary of the latest approaches that are based on recurrent models for sequential data processing. A complete taxonomy over the latest trends in architectural and algorithmic solutions is reported and discussed, guiding researchers in this appealing research field. The emerging picture suggests that there is room for thinking of novel routes, constituted by learning algorithms which depart from the standard Backpropagation Through Time, towards a more realistic scenario where patterns are effectively processed online, leveraging local-forward computations, opening to further research on this topic.
@@ 1 Introduction
Human cognition is inherently intertwined with the seamless processing of sequential patterns [65, 270, 48]. Sequential data is clearly ubiquitous in the context of perception: from the flow of natural language during conversations to the sequence of cues processed in visual perception, and more. Mimicking the human ability to comprehend and learn from sequences has long been an aspiration within the realm of Artificial Intelligence (AI), with the ultimate goal of mirroring the human capacity to retain long-term insights from past experiences while remaining attuned to upcoming information [273, 169]. The relevance of processing sequential data is evidenced by the increasing amount of tasks tackled with Deep Learning for scientific and industrial applications, such as conversational AI [28], natural language understanding [294], video representation learning and processing [263], lifelong and continual learning [273, 62], time-series analysis [21], temporal graphs [307], and many others [11, 128].
@@
Early attempts to deal with sequential data are dated back to the dawn of Machine Learning [187, 247, 108, 65, 220], and they focused on the proposal of novel architectural and algorithmic solutions that departed from popular approaches for static data, and more apt to the context of processing sequences. Indeed, Recurrent Neural Networks (RNNs) go beyond feed-forward models for static data thanks to the way hidden states are structured, that work as the memory of the network and are conditioned by previous-time self-connections [65, 253]. Given the novelty of such architectural topologies, several ad-hoc learning rules were proposed [303, 302], among which the Backpropagation Through Time (BPTT) [299] algorithm emerged and became prominent. However, BPTT necessitates to recursively unfold the network over the processed sequences, storing intermediate computations over the whole input sequence, and virtually obtaining a “deep” feed-forward computational graph on which standard back-propagation of errors is applied [209].
@@
The limits and drawbacks of such solution emerge when the goal is the one of emulating abilities which are typical of human cognition, such as real-time/online sequence processing. Indeed, the unfolded networks represent very deep pathways to be traversed by error information. As a result, RNNs trained by BPTT suffer from vanishing gradients [18, 273] making credit assignment much more difficult [208], i.e., the task of computing the impact on the overall error of the individual units in the network. Early solutions often necessitated compromises to make learning affordable, simplifying the learning paradigm by reducing the context window (e.g., Truncated BPTT [302]) or assuming the availability of complete sequences in advance. Yet, the aforementioned issues and these simplifications constrained the ability to capture intricate long-term dependencies and correlations that characterize many real-world sequential datasets [88].
@@
Figure 1: Left: Since the advent of Transformers [291], the neural network community promoted such architectures as alternatives to recurrent (RNNs) and convolutional (CNNs) models for sequence processing, i.e., “attention is all you need”, thus they are represented by disjoint boxes in the picture, with the main features characterizing each of them listed at the top. Right: In the current generation of neural models, described in this survey, attention “married” recurrence, generating a significant intersection between the different “worlds”. RNNs are state-space models, and in the same few years we observed the growth of several variants of (deep) State-Space Models, where linearity (yellow set) is one of the main keys behind their success (together with element-wise recurrence and gating functions), and convolutional layers are now parts of various architectures. Interestingly, linearity intersects all the three families of models (RNNs, Transformers, Deep State-Space Models—the latter are almost fully based on linear models). The increasing interest in lifelong-Learning and applications running on limited-computation devices and with privacy constraints, such as instances of edge computing, is shifting the interest toward online learning (previously a prerogative of stateful/RNNs models). The white area is the meeting point of all the important features listed on top, and where there might be more room for novel research activities.
@@
However, the ultimate goal of mimicking the human capacity to learn over sequences in real-time inspired several researchers. Amongst others, Williams and Peng proposed “an on-line algorithm, designed to be used to train a network while it runs; no manual state resets or segmentations of the training stream is required” [302], as long as other best known variants (Real Time Recurrent Learning, RTRL [303]) that however suffer from scalability issues [122]. Architectural solutions were investigated, for instance through the introduction of gating mechanisms (Long Short Term Memories, LSTMs [105]) that partially improves the control over vanishing gradients and the error flow. Remarkably, the original work on gating in LSTMs [76] was devised as a remedy to the fact that “a continual input stream eventually may cause the internal values of cells to grow without bound”. Still, none of these solutions was capable to overcome the challenges posed by limited computational resources and the quest for efficiency in sequence processing in RNNs, ending up in the inability to process sequences longer than few thousands steps [293, 157]. Going beyond instances of RNNs, still in the context of neural networks, the scientific literature includes approaches to long sequence processing which are based on Convolutional Neural Networks (CNNs)—see [101] and references therein. CNNs promoted the capability of parallelizing inference/learning over the temporal dimension, when the entire sequence is available in advance [13]. The local nature of the convolution is triggered by the design choice of using filters that span on a limited time range around each time instant. However, dealing with small filters hinder the ability to capture very long-term dependencies, a role that is delegated to the effect of stacked convolutional layers.
@@
The ubiquity of Transformers [291] in the last years has lately overtaken the sequence processing scenario due to several advantages with respect to the other existing architectures. In fact, the training procedure is parallelizable on modern hardware when the full sequence is available [318]. The self-attention mechanism introduces several advantages, for instance the ability of handling both local and long-range relations [282] completely avoiding the vanishing gradient issue, thanks to the direct connection of any token pairs in the sequence. However, the quadratic complexity characterizing self-attention yields a significant computational cost and memory demands, particularly pronounced when handling long input sequences and strongly limiting in case of on-the-edge devices with small computational resources. These issues have instigated a profusion of research endeavors with the goal of refining the scalability attributes of Transformers, often entailing trade-offs with certain traits that underlie their efficiency [295, 54]. Several of these approaches are inspired by intuitions coming from RNNs, that are fast-inference machines that scale linearly with the sequence length [222].
@@
In the same years during which Transformers spread over the Machine Learning universe, and inspired by approaches devised for continuous-time recurrent neural networks [293, 86], an alternative path to handle long-range sequences emerged [88], based on state-space models, as RNNs. A particular instance of state-space models was more strongly considered by scientists, i.e., linear State-Space Models (SSM), which promoted the idea of avoiding the direct application of non-linearities in the state-update rule. When discretized with ad-hoc integrators, continuos time SSMs were used as additional modules to gated RNNs [293, 86]. This intuitions inspired a plethora of works aiming at injecting linear SSMs into Deep Architectures [90, 95, 271], frequently considering diagonal recurrent matrices. More recently, the authors of [213] proposed a recurrent model that bridges RNNs and the intuitions behind the aforementioned categories of SSMs, referred to as Linear Recurrent Units (LRUs). Basically, they showed that appropriately parametrizing and structuring RNNs (linear and diagonal recurrence, specific initializations, etc.) the advantages of Deep-SSMs can be fully exploited also in standard RNNs. When paired with appropriate gating functions and very structured deep networks, these models can reach state-of-the art results in language modeling [57]. Of course, while it is useful to talk about Deep-SSMs and RNNs in a separated manner to better emphasize the recent literature on linear models (frequently based on diagonal recurrent matrices and additional neural layers on top of them), we remark that also RNNs are indeed state-space models. We summarize in Table 1 the aforementioned ingredients constituting the novel components of current solutions, that we will deepen in the remainder of the paper. While previous approaches did not consider many of this architectural components (linear and element-wise recurrence, gating mechanism), they have proven fundamental to be capable to deal with the increasing demanding long-term sequences. When the sequence length is taken to the extreme, i.e. dealing with potentially infinite sequences, the current learning paradigm based upon BPTT should be re-discussed by considering online learning scenarios.
@@
Current Gen Sec.3: Transformers embracing Recurrence -
@@
Motivation and Scope.   In the present era, a confluence of factors, including the rise of large-scale language models [28] and the need of optimizing their performances on both training and inference, resulted in a resurgence of interest on recurrent architectures [213, 222] and their optimization schemes, revolving into a new chapter in the narrative of sequence processing, going beyond the “attention is all you need” conjecture [291] (Figure 1). This survey is a comprehensive exploration of such a chapter, encompassing the latest architectural solutions inspired by recurrent computations, ranging from Transformers with extensive context to the rekindling of interest in linear recurrent networks, promoted by multiple variants of deep state-space models. Meanwhile, innovative learning algorithms that challenge the conventional Backpropagation Through Time (BPTT) have emerged. These algorithms embody a departure from the conventional approach, embracing the practical need for online sequence processing [184]. They leverage localized and forward-facing strategies to navigate the nuances of real-time pattern recognition [122]. The primary goal of this survey is to offer an exhaustive taxonomy of contemporary trends in both architectural design and algorithmic innovation for sequence processing with recurrent architectures. By shedding light on the cutting-edge techniques, we aim to provide researchers with a road map that navigates the dynamic landscape of sequence processing, guiding them in their pursuit of effective solutions. As the AI community stands on the threshold of unprecedented opportunities catalyzed by the advent of large language models, we invite researchers to embark on this expedition through the intricate realm of sequence processing. This survey also opens toward novel research opportunities, that lie at the intersection of online learning from sequential data and lifelong learning [177]. Figure 1 summarizes the big picture of this paper, emphasizing the “marriage” between RNNs, Transformers, and the evolution of state-space models based on linear units. In a nutshell, stateful models are not only back in the context of learning from sequential data, but they also bring together approaches that were previously strongly kept separated by the scientific community. Novel opportunities might lie at the intersection with online lifelong learning.
@@
Paper Overview.   Figure 2 describes the way this paper is organized. In each section which surveys multiple categories of models, we will further provide a picture to summarize its contents and main methods. We start this survey by describing in Section 2 recurrent models for sequence processing and introducing the notation we will use in the remainder of this manuscript. Moreover, we will include references to previous efforts in surveying sequence processing based on recurrent neural networks. We will formalize our derivations using a notation that will be exploited across apparently different models, favouring comparisons among them. Section 3 surveys the latest findings towards the optimization of Transformers for long sequence processing, focusing on methods that search for optimized self-attention mechanisms inspired by recurrent approaches. Section 4 provides a thorough description of the emerging field of deep state-space models. In Section 6 we focus on learning dynamics, describing the recent learning algorithms that challenge the conventional Backpropagation Through Time (BPTT), searching for local and forward-facing methods, aiming at emulating the human ability to process sequences in an online manner. Then, in Section 7 we survey the most recent benchmarks that are used to test models in grasping long-term dependencies. Finally, in Section 8 we analyze some open problems and issues that have been recently pointed out in the context of the described works, as long as highlighting possible future avenues for research, and we draw our conclusions in Section 9.
@@ 2 Recurrent Models in Sequence Processing
In several real-world scenarios, data is organized in a sequential manner, where ordering does matter and which requires appropriate computational models. It is the case of natural language text, speech, vision, time series, and others. Leveraging the temporal relations and ordering of patterns can help in discovering many additional cues in data [65]. Let us describe a sequence of $L$ patterns with the notation
@@
where $L$ is a positive integer that represents the sequence length and $u_{t}\in\mathbb{R}^{{d_{\text{in}}}}$, $t=1,\dots,L$, is the $t$-th pattern. [^1] This holds both when $u_{t}$ is considered to be the external input of a computational model or the result of intermediate computations in a multi-layer network. Notice that while $t$ is commonly intended to be as the index of a pattern in the sequence, in many settings (especially in the context of perception or of time-series) we also have the use of the time at which the sample was provided or, equivalently, of the length of the time interval that passed between consecutive patterns. We also remark that, in principle, $L$ could also be infinite. In this paper, to keep the notation simple, we use $t$ both as index of pattern/step and as time variable, depending on the context.
@@
Stateless Models.   The most straightforward approach for handling sequences, which have been considered since before recurrent architectures became ubiquitous, is to explicitly inject time by means of a spatial representation (as discussed in [65], for example). The order of events in a sequence is simply considered to be an additional dimension of the pattern, which can be processed in a single shot. This also allows to designed models that process in parallel the information at the different time instants, i.e., without having to wait for computations on past data to finish. There exists several works in the late eighties described this route [66, 96], which has been also followed by more recent works based on convolutions [88] and Transformers [291]. The former virtually slides fixed-length filters over the input sequence [75, 271], the latter leverages positional encodings, in both the cases parallelizing computations over time [61]. Overall, these models are stateless, in the sense that they do not try to build and progressively update a representation of the sequence observed so far. As a consequence, they require the whole sequence to be available and/or to reconsider it all if it gets updated, thus not being suited, for example, for continuous online-learning [79].
@@
Recurrent Models.   An alternative approach consists in focusing on stateful models that specifically embrace time in their computational scheme, developing a form of memory which gets updated over time (i.e., the state). This is achieved by introducing feedback connections, yielding instances of what is commonly referred to as Recurrent Models. Processing a sequence in Recurrent Models consists in handling one pattern at-a-time, following the original order of the data. Internal units are used to model the so-called state or context, and, due to the feedback connections, they are responsible of fusing information from the current patterns and information from the previous time step, in order to update the state representation. As a result, the state effectively encodes the temporal characteristics of the sequential input observed so far [65]. Several seminal works investigated this very natural direction [247, 298]. Jordan [127] introduced a network with one-to-one connections from output units to state units, as shown in Figure 3-left. Feedback connections enable the hidden units to access the prior outputs of the network, thereby shaping subsequent behaviors. This property grants the network a form of memory, facilitating the association of static patterns (referred to as “Plans”) with serially ordered output patterns (sequences of “Actions”). Elman [65] proposed the best known instance of Recurrent Model, where state units interact exclusively with internal nodes of the network, and not with the external world (i.e., state units are fed with the hidden information of the previous step, instead of what comes from the output layer), as shown in Figure 3-right. There exists an important connection between stateful Recurrent Model and the generic notion of state-space model, which has been used in many fields of engineering [145]. This notion is common in control theory to model a dynamical system, and it is generic enough to cover a large number of possible implementations, covering linear and non-linear systems, being them time variant or invariant. However, in the context of the majority of literature in Machine Learning, it is rarely mentioned together with Recurrent Models. State-space models define the temporal evolution of state variables by first-order differential equations or difference equations. The state changes over time in function of its value at a given instant and on the currently provided external input. This definition intrinsically covers the typical feedback connection of Recurrent Models, which can be considered instances of state-space models, as we will formalize in what follows.
@@
Figure 3: Jordan (left) and Elman (right) implementations of recurrent models. $A$, $B$, $C$ represent linear projections, and input (triangles), hidden (circles-blue boxed), output (orange-boxed) units are emphasized ($\sigma$, and $\sigma_{y}$ are activation functions).
@@
Recurrent Neural Networks.   Elman’s architecture [65] was the pillar to the foundations of the widespread notion of Recurrent Neural Network (RNN), summarizing a neural model with a special hidden-layer including lookback connections, that we refer to as RNN layer. Given an input sequence $(u_{1},u_{2},\dots,u_{L})$, an RNN layer with ${d_{\text{s}}}$-dimensional hidden state $x_{k}$ computes a sequence of ${d_{y}}$-dimensional outputs $(y_{1},y_{2},\dots,y_{L})$ through a recurrent model,
@@
Moving the term $-x(t)$ to the left-hand side of the first equation yields an evident connection to Eq. (2). This formulation is well-suited to trace another important link with the already introduced notion of state-space model. The general form of a state-space model is actually close to the one of Eq. (3), when also including the direct dependence on time $t$ in both the equations, i.e., $\dot{x}(t)=f(x(t),u(t),t)$ and $y(t)=h(x(t),u(t),t)$, being $f$ and $h$ two generic functions. The discrete counterpart of the first equation, when considering time invariant systems, intrinsically yields the classic feedback of Recurrent Models.

### v2

@@ Abstract
Effectively learning from sequential data is a longstanding goal of Artificial Intelligence, especially in the case of long sequences. From the dawn of Machine Learning, several researchers have pursued algorithms and architectures capable of processing sequences of patterns, retaining information about past inputs while still leveraging future data, without losing precious long-term dependencies and correlations. While such an ultimate goal is inspired by the human hallmark of continuous real-time processing of sensory information, several solutions have simplified the learning paradigm by artificially limiting the processed context or dealing with sequences of limited length, given in advance. These solutions were further emphasized by the ubiquity of Transformers, which initially overshadowed the role of Recurrent Neural Nets. However, recurrent networks are currently experiencing a strong recent revival due to the growing popularity of (deep) State-Space models and novel instances of large-context Transformers, which are both based on recurrent computations that aim to go beyond several limits of currently ubiquitous technologies. The fast development of Large Language Models has renewed the interest in efficient solutions to process data over time. This survey provides an in-depth summary of the latest approaches that are based on recurrent models for sequential data processing. A complete taxonomy of recent trends in architectural and algorithmic solutions is reported and discussed, guiding researchers in this appealing research field. The emerging picture suggests that there is room for exploring novel routes, constituted by learning algorithms that depart from the standard Backpropagation Through Time, towards a more realistic scenario where patterns are effectively processed online, leveraging local-forward computations, and opening new directions for research on this topic.
@@ 1 Introduction
Human cognition is inherently intertwined with the seamless processing of sequential patterns [1, 2, 3]. Sequential data are ubiquitous in the context of perception: from the flow of natural language during conversations to the sequence of cues processed in visual perception, and more. Mimicking the human ability to comprehend and learn from sequences has long been an aspiration within the realm of Artificial Intelligence (AI), with the ultimate goal of mirroring the human capacity to retain long-term insights from past experiences while remaining attuned to upcoming information [4, 5]. The relevance of processing sequential data is evidenced by the increasing amount of tasks tackled with Deep Learning for scientific and industrial applications, such as conversational AI [6], natural language understanding [7], video representation learning and processing [8], lifelong and continual learning [4, 9], time-series analysis [10], temporal graphs [11], and various other domains[12, 13].
@@
Early attempts to deal with sequential data date back to the dawn of Machine Learning [14, 15, 16, 1, 17], and they focused on the proposal of novel architectural and algorithmic solutions that departed from popular approaches designed for static data, and were better suited to the context of processing sequences. Indeed, Recurrent Neural Networks (RNNs) go beyond feed-forward models for static data thanks to the way hidden states are structured, serving as a form of memory and being influenced by previous-time self-connections [1, 18]. Given the novelty of such architectural topologies, several ad-hoc learning rules were proposed [19, 20], among which the Backpropagation Through Time (BPTT) [21] algorithm emerged and became prominent. However, BPTT requires the recursive unfolding of the network over the processed sequences, storing intermediate computations across the whole input sequence, and virtually obtaining a “deep” feed-forward computational graph on which standard back-propagation of errors is applied [22].
@@
The limits and drawbacks of this solution become evident when the goal is to emulate abilities which are typical of human cognition, such as real-time/online sequence processing. Indeed, the unfolded networks represent very deep pathways to be traversed by error information. As a result, RNNs trained by BPTT suffer from vanishing gradients [23, 4] making credit assignment much more difficult [24], i.e., the task of computing the impact on the overall error of the individual units in the network. Early solutions often necessitated compromises to make learning affordable, simplifying the learning paradigm by reducing the context window (e.g., Truncated BPTT [20]) or assuming the availability of complete sequences in advance. Yet, the aforementioned issues and these simplifications constrained the ability to capture intricate long-term dependencies and correlations that characterize many real-world sequential datasets [25].
@@
Figure 1: Left: Since the advent of Transformers [26], the neural network community has promoted these architectures as alternatives to recurrent (RNNs) and convolutional (CNNs) models for sequence processing, i.e., “attention is all you need”, thus they are represented by disjoint boxes in the picture, with the main features characterizing each of them listed at the top. Right: In the current generation of neural models, described in this survey, attention has “married” recurrence, generating a significant intersection between the different “worlds”. RNNs are state-space models, and in recent years we have observed the emergence of several variants of (deep) State-Space Models, where linearity (yellow set) is one of the main keys behind their success (together with element-wise recurrence and gating functions), and convolutional layers are now parts of various architectures. Interestingly, linearity intersects all the three families of models (RNNs, Transformers, Deep State-Space Models—the latter are almost fully based on linear models). The increasing interest in lifelong-Learning and applications running on limited-computation devices and with privacy constraints, such as instances of edge computing, is shifting the interest toward online learning (previously a prerogative of stateful/RNNs models). The white area is the meeting point of all the important features listed on top, and where there might be more room for novel research activities.
@@
However, the ultimate goal of mimicking the human capacity to learn over sequences in real-time has inspired several researchers. Amongst others, Williams and Peng proposed“an on-line algorithm, designed to be used to train a network while it runs; no manual state resets or segmentations of the training stream are required” [20], as well as other well-known variants such as Real Time Recurrent Learning (RTRL) [19], which, however, suffer from scalability issues [27]. Architectural solutions were investigated, for instance through the introduction of gating mechanisms (Long Short Term Memories, LSTMs [28]) that partially improve the control over vanishing gradients and the error flow. Remarkably, the original work on gating in LSTMs [29] was devised as a remedy to the fact that “a continual input stream eventually may cause the internal values of cells to grow without bound”. Still, none of these solutions was capable of overcoming the challenges posed by limited computational resources and the quest for efficiency in sequence processing in RNNs, ultimately resulting in their inability to process sequences longer than a few thousand steps [30, 31]. Going beyond instances of RNNs, still in the context of neural networks, the scientific literature includes approaches to long sequence processing that are based on Convolutional Neural Networks (CNNs)—see [32] and references therein. CNNs enable the capability of parallelizing inference/learning over the temporal dimension when the entire sequence is available in advance [33]. The local nature of the convolution results from the design choice of using filters that span a limited time range around each time instant. However, dealing with small filters hinder s the ability to capture very long-term dependencies, a role that is instead fulfilled by stacking multiple convolutional layers.
@@
The ubiquity of Transformers [26] in recent years has dominated the sequence processing scenario due to several advantages over other existing architectures. In fact, the training procedure is parallelizable on modern hardware when the full sequence is available [34]. The self-attention mechanism introduces several advantages, such as the ability to handle both local and long-range relations [35], while completely avoiding the vanishing gradient issue, thanks to the direct connection of any token pairs in the sequence. However, the quadratic complexity characterizing self-attention yields a significant computational cost and memory demand, particularly pronounced when handling long input sequences, which becomes a strong limitation in the case of on-the-edge devices with limited computational resources. These issues have instigated a profusion of research endeavors aimed at improving the scalability attributes of Transformers, often entailing trade-offs with certain traits that underlie their efficiency [36, 37]. Several of these approaches are inspired by insights derived from RNNs, that are fast-inference machines that scale linearly with the sequence length [38].
@@
In parallel to these developments, and during the same years in which Transformers were gaining widespread adoption, an alternative and independent research direction emerged, based upon the usage of State-Space Models (SSMs) for sequence modeling. Inspired by methods developed for continuous-time recurrent neural networks [30, 39], this line of work proposed SSMs as a principled way to handle long-range dependencies while preserving efficiency. In particular, a subclass of linear SSMs gained traction, promoting the idea of removing nonlinearities from the state-update rule to enable efficient computations [25]. When discretized with ad-hoc integrators, continuos time SSMs were used as additional modules to gated RNNs [30, 39]. This intuitions inspired a plethora of works aiming at injecting linear SSMs into Deep Architectures [40, 41, 42], frequently considering diagonal recurrent matrices and further blurring the line between these model classes. More recently, the authors of [43] proposed a recurrent model that bridges RNNs and the intuitions behind the aforementioned categories of SSMs, referred to as Linear Recurrent Units (LRUs). Basically, they showed that appropriately parametrizing and structuring RNNs (linear and diagonal recurrence, specific initializations, etc.) the advantages of Deep-SSMs can be fully exploited also in standard RNNs. When paired with appropriate gating functions and very structured deep networks, these models can reach state-of-the art results in language modeling [44]. While it is useful to talk about Deep-SSMs and RNNs in a separate manner to better emphasize the recent literature on linear models (frequently based on diagonal recurrent matrices and additional neural layers on top of them), we note that RNNs are also instances of state-space models. We summarize in Table 1 the aforementioned ingredients constituting the novel components of current solutions, which we will explore in greater detail in the remainder of the paper. While previous approaches did not consider many of these architectural components—such as linear and element-wise recurrence and gating mechanisms— they have proven fundamental for effectively managing increasingly demanding long-term sequences. When the sequence length is taken to the extreme, i.e., dealing with potentially infinite sequences, the current learning paradigm based upon BPTT must be reconsidered by considering online learning scenarios.
@@
Current Gen
Sec.3: Transformers embracing Recurrence -
@@
Motivation and Scope.   In the present era, a confluence of factors, including the rise of large-scale language models [6] and the need to optimize their performance during both training and inference, has resulted in a resurgence of interest in recurrent architectures [43, 38] and their optimization schemes, marking a new chapter in the narrative of sequence processing, going beyond the “attention is all you need” conjecture [26] (Figure 1). This survey presents a comprehensive exploration of this recent evolution, covering the latest architectural solutions inspired by recurrent computations, ranging from Transformers with extensive context to the rekindling of interest in linear recurrent networks, driven by multiple variants of deep state-space models. Meanwhile, innovative learning algorithms that challenge the conventional Backpropagation Through Time (BPTT) have emerged. These algorithms represent a significant shift from the traditional methodology, addressing the practical need for online sequence processing [45]. They leverage localized and forward-facing strategies to navigate the nuances of real-time pattern recognition [27]. The primary goal of this survey is to offer an exhaustive taxonomy of contemporary trends in both architectural design and algorithmic innovation for sequence processing with recurrent architectures. By shedding light on the cutting-edge techniques, we aim to provide researchers with a comprehensive guide through the dynamic landscape of sequence processing, guiding them in their pursuit of effective solutions. As the AI community stands on the threshold of unprecedented opportunities catalyzed by the advent of large language models, we invite researchers to embark on this journey into the intricate realm of sequence processing. This survey also highlights emerging research directions that lie at the intersection of online learning from sequential data and lifelong learning [46]. Figure 1 summarizes the big picture of this paper, emphasizing the“marriage” between RNNs, Transformers, and the evolution of state-space models based on linear units. In a nutshell, stateful models are not only back in the context of learning from sequential data, but they also bring together approaches that were previously considered distinct by the scientific community. New research opportunities may emerge at the intersection with online lifelong learning.
@@
Paper Overview.   Figure 2 describes the way this paper is organized. In each section that surveys multiple categories of models, we will further provide a picture to summarize its contents and main methods. Moreover, we include references to previous efforts in surveying sequence processing based on recurrent neural networks. In Table 2, we report the main mathematical symbols used throughout the paper, together with their descriptions and dimensionalities. We start this survey by describing, in Section 2, recurrent models for sequence processing and introducing the notation we will use in the remainder of this manuscript. We will formalize our derivations using a notation that is applied consistently across apparently different models, to facilitate comparisons among them. Section 3 surveys the latest findings regarding the optimization of Transformers for long sequence processing, focusing on methods that aim to improve self-attention mechanisms inspired by recurrent approaches. Section 4 provides a thorough description of the emerging field of deep state-space models. In Section 6, we focus on learning dynamics, describing the recent learning algorithms that challenge the conventional Backpropagation Through Time (BPTT), in pursuit of local and forward-facing methods designed to emulate the human ability to process sequences in an online manner. Then, in Section 7, we survey the most recent benchmarks used to evaluate how well models capture long-term dependencies. Finally, in Section 8, we analyze some open problems and issues that have been recently pointed out in the context of the described works, as well as highlight possible future avenues for research, and we draw our conclusions in Section 9.
Table 2: Summary of the main mathematical symbols used throughout the paper, including RNNs, Transformers, and Deep State-Space Models, along with their description and dimensionality.
Symbol Description Dimension
$L$ Sequence length -
$x$ Hidden state $d_{s}$
$u$ Input vector $d_{\text{in}}$
$y$ Output vector $d_{y}$
AA State transition matrix $d_{s}\times d_{s}$
$B$ Input matrix $d_{s}\times d_{\text{in}}$
$C$ Output matrix $d_{y}\times d_{s}$
$D$ Feedthrough matrix $d_{y}\times d_{\text{in}}$
$\theta$ Learnable parameters -
$\sigma$ Activation function -
$q,k$ Query, key, value vectors $d_{k}$
$v$ Value vectors $d_{y}$
$U$ Stacked input (sequence) $L\times d_{\text{in}}$
$Q,K$ Stacked query/key (sequence) $L\times d_{k}$
$V$ Stacked value (sequence) $L\times d_{y}$
$W_{q},W_{k}$ Linear query/key projection matrices $d_{k}\times d_{\text{in}}$
$W_{v}$ Linear value projection matrices $d_{y}\times d_{\text{in}}$
@@ 2 Recurrent Models in Sequence Processing
In several real-world scenarios, data is organized in a sequential manner, where ordering matters and requires appropriate computational models. This is the case of natural language text, speech, vision, time series, and others. Leveraging the temporal relations and ordering of patterns can help in discovering many additional cues in data [1]. Let us describe a sequence of $L$ patterns with the notation
@@
where $L$ is a positive integer that represents the sequence length and $u_{t}\in\mathbb{R}^{{d_{\text{in}}}}$, $t=1,\dots,L$, is the $t$-th pattern. [^1] This holds both when $u_{t}$ is considered to be the external input of a computational model or the result of intermediate computations in a multi-layer network. Note that while $t$ is commonly interpreted as the index of a pattern in the sequence, in many settings (especially in the context of perception or time-series) we also have the use of the time at which the sample was provided or, equivalently, of the length of the time interval that passed between consecutive patterns. We also remark that, in principle, $L$ could be infinite. In this paper, to keep the notation simple, we use $t$ both as the index of a pattern or step and as the time variable, depending on the context. [^2] In the following, the symbol $t$ will be written as a subscript when referring to a discrete-time index, for example $u_{t}$, and enclosed in parentheses when representing a continuous-time variable, for example $u(t)$.
@@
Stateless Models.   The most straightforward approach for handling sequences, which have been considered since before recurrent architectures became ubiquitous, is to explicitly inject time by means of a spatial representation (as discussed in [1], for example). The order of events in a sequence is simply considered to be an additional dimension of the pattern, which can be processed in a single shot. This also allows to design models that process in parallel the information at the different time instants, i.e., without having to wait for computations on past data to finish. There exists several works in the late eighties described by this route [47, 48], which has been also followed by more recent works based on convolutions [25] and Transformers [26]. The former virtually slides fixed-length filters over the input sequence [49, 42], the latter leverages positional encodings, in both cases parallelizing computations over time [50]. Overall, these models are stateless, in the sense that they do not try to build and progressively update an internal representation of the sequence observed so far. As a consequence, they require the whole sequence to be available and/or to reconsider it all if it gets updated, thus not being suited, for example, for continuous online-learning [51].
@@
Recurrent Models.   An alternative approach consists in focusing on stateful models that specifically embrace time in their computational scheme, developing a form of memory that gets updated over time (i.e., the state). This is achieved by introducing feedback connections, yielding instances of what are commonly referred as Recurrent Models. Processing a sequence in Recurrent Models consists of processing one pattern at-a-time, following the original order of the data. Internal units are used to model the so-called state or context, and, due to the feedback connections, they are responsible for integrating information from the current patterns and information from the previous time step, in order to update the state representation. As a result, the state effectively encodes the temporal characteristics of the sequential input observed so far [1]. Several seminal works investigated this very natural direction [15, 52]. Jordan [53] introduced a network with one-to-one connections from output units to state units, as shown in Figure 3-left. Feedback connections enable the hidden units to access the prior outputs of the network, thereby shaping subsequent behaviors. This property grants the network a form of memory, facilitating the association of static patterns (referred to as “Plans”) with serially ordered output patterns (sequences of “Actions”). Elman [1] proposed the best known instance of a Recurrent Model, where state units interact exclusively with internal nodes of the network, and not with the external world (i.e., state units are fed with the hidden information of the previous step, instead of what comes from the output layer), as shown in Figure 3-right. There exists an important connection between stateful Recurrent Models and the generic notion of state-space model, which has been used in many fields of engineering [54]. This notion is common in control theory to model a dynamical system, and it is generic enough to cover a large number of possible implementations, including linear and non-linear systems, whether time-varying or time-invariant. However, in the context of the majority of the literature in Machine Learning, it is rarely mentioned together with Recurrent Models. Koopman theory offers a principled alternative to RNNs for modeling nonlinear dynamical systems by embedding their evolution into a linear function space via the Koopman operator. This perspective enables more stable and interpretable models, as testified by works that incorporate Koopman theory into deep learning through physics-informed autoencoders that outperform traditional RNNs [55]. See Section 6 for further discussion on the connections between RNNs and dynamical systems.
State-space models define the temporal evolution of state variables by first-order differential equations or difference equations. The state changes over time as a function of its value at a given instant and the currently provided external input. This definition intrinsically covers the typical feedback connection of Recurrent Models, which can be considered instances of state-space models, as we will formalize in what follows.
@@
Figure 3: Jordan (left) and Elman (right) implementations of recurrent models. AA, $B$, $C$ represent linear projections, and input (triangles), hidden (circles-blue boxed), output (orange-boxed) units are emphasized ($\sigma$, and $\sigma_{y}$ are activation functions).
@@
Recurrent Neural Networks.   Elman’s architecture [1] was the pillar to the foundations of the widespread notion of Recurrent Neural Network (RNN), summarizing a neural model with a special hidden-layer including lookback connections, that we refer to as RNN layer. Given an input sequence $(u_{1},u_{2},\dots,u_{L})$, an RNN layer with $d_{\text{s}}$}-dimensional hidden state $x_{t}$ computes a sequence of $d_{y}$}-dimensional outputs $(y_{1},y_{2},\dots,y_{L})$ through a recurrent model,
@@
Moving the term x(t)-x(t) to the left-hand side of the first equation yields an evident connection to Eq. (2). This formulation is well-suited to trace another important link with the already introduced notion of state-space model. The general form of a state-space model is actually close to the one of Eq. (3), when also including the direct dependence on time $t$ in both the equations, i.e., $\dot{x}(t)=f(x(t),u(t),t)$ and $y(t)=h(x(t),u(t),t)$, being $f$ and $h$ two generic functions. The discrete counterpart of the first equation, when considering time invariant systems, intrinsically yields the classic feedback of Recurrent Models.